OpenStudy (anonymous):

What is the radius of a hydrogen atom whose electrom moves at 2.4 x 10^5 m/s?

OpenStudy (anonymous):

oh, nice! You will have to use centripetal force, and equate it to the coloumb force..? \[ \frac{q_{1}q_{2}}{4 \pi \epsilon_{0} r^{2}} = \frac{m_{e}v^{2}}{r} \] all of the letters are constants you can look up, except for \(v \) which you know, and \( r \) which you solve for.